Is the order pair (3,8) a solution to the system of equations 2x+y=14. 3x-2y=7
Why or why not show your work please Answer by harpazo(655) 
 In the point (3,8), x = 3 and y = 8.
When I replace x with 3 and y with 8 in both equations given and do the math, the answer is not the same on both sides of the linear equations. So, the point (3,8) is not a solution to this system of linear equations.
The Prove
Let x = 3
Let y = 8
2x+y=14
2(3) + 8 = 14
6 + 8 = 14
14 = 14
It works for this equation. Let us now do the same with the other equation.
3x-2y=7
3(3) - 2(8) = 7
9 - 16 = 7
-7 DOES NOT = 7
Since the point (3,8) does not produce the same answer on both sides of each equation, the given point is not a solution to this system of linear equations.
Understand?